OpenStudy (anonymous):

4 sinx sin(pi/3+x) sin(2 pi/3+x) = sin3x ? ?

OpenStudy (solomonzelman):

\[π = 180\] So,

OpenStudy (solomonzelman):

Lets convert radians to degrees. 4 sinx sin(pi/3+x) sin(2 pi/3+x) = sin3x 4 sinx sin(60+x) sin(120+x) = sin3x sin(60 + x) = sin(60)cos(x) + cos(60)sin(x) sin(120 + x) = sin(120)cos(x) + cos(120)sin(x) So we have so far, (I'm assuming it's 4sinx times...since there is no plus symbol) 4 sinx (sin(120)cos(x) + cos(120)sin(x) + sin(60)cos(x) + cos(60)sin(x))= sin3x) = sin3x sin(120)cos(x) + cos(120)sin(x) + sin(60)cos(x) + cos(60)sin(x))= sin3x= sin3x / 4sinx

OpenStudy (solomonzelman):

\[Sin120=Sin60=\frac{\sqrt{3}}{2}\]\[Cos120=-\frac{1}{2}\]\[Cos60=\frac{1}{2}\]

OpenStudy (solomonzelman):

In my second comment, ignore at the end where it says "=sin(3x)" the first time